OpenStudy (anonymous):

Use algebra to prove the identity: sin t/1-cos t = 1+cos t/ sin t

OpenStudy (anonymous):

sin t/1-cost=sint/(1-cost)*[(1+cos)/(1+cost)] = sint(1+cost)/(1-cos^2 t) = sint(1+cost)/(sin^2 t) = (1+cost)/sint